OpenStudy (anonymous):

You are driving to the grocery store at 20 . You are 110 from an intersection when the traffic light turns red. Assume that your reaction time is 0.50 and that your car brakes with constant acceleration. How far are you from the intersection when you begin to apply the brakes?

OpenStudy (jamesj):

the units are 20 m/s (km/hr? m/hr?) 110 ? 0.5 seconds?

OpenStudy (anonymous):

meters and seconds

OpenStudy (jamesj):

Ok, so far have you travelled in 0.5 seconds?

OpenStudy (jamesj):

velocity, v = 20 m/s time, t = 0.5 displacement, d. Now by definition of velocity v = d/t. Hence d = vt = ....what?

OpenStudy (anonymous):

40 m

OpenStudy (anonymous):

no 10 m right?

OpenStudy (jamesj):

Distance = velocity * time d=vt=(20 m/s)(0.5 s)= 10 m So 10 meters, yes

OpenStudy (jamesj):

Hence what's the answer to the question?

OpenStudy (jamesj):

At 110 meters away when you see the red light. Half a second goes by, you travel 10 m and then start to break. How far away from the intersection are you now?

OpenStudy (anonymous):

100 m

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

thanks, you're awsome!

OpenStudy (jamesj):

yes